rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_green
icon_orange
icon_red
icon_green
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#4946 — FS#8876 — Optical maintenance LDN <> AMS

Attached to Project— Network
Maintenance
Whole Network
CLOSED
100%
We will be carrying out maintenance works on the fibre optic between Londres and Amsterdam to resolve imbalance and weakening issues between the two cables on the Wickford <> London stretch.

The Wickford <> Londres stretch will be replaced by a new pair of fibres with a optimised line.
Date:  Thursday, 27 June 2013, 15:31PM
Reason for closing:  Done
Comment by OVH - Thursday, 27 June 2013, 11:51AM

We are starting by changing the patch at Wickford.


Comment by OVH - Thursday, 27 June 2013, 12:46PM

The patch was modified at Wickford, now we are doing the same on the London side.


Comment by OVH - Thursday, 27 June 2013, 15:30PM

The photonic link is back up. We have gained 10dB int the Wickford towards London direction and 8dB from London to Wickford.

Part optimisation of Raman amps is in progress.


Comment by OVH - Thursday, 27 June 2013, 15:31PM

Optimisation is finished. The optic link is operational again.